OpenStudy (anonymous):

1) Find exact solution a) Using identity: sin x + Sqrt(3) cos x = 1, where 0 =< x =< 2pi b) Using tan 2 thetha , thetha [-pi,pi] : (2sin thetha)/cos thetha = tan^2 thetha - 1 c) Cos 4x = -2cos^2 x

OpenStudy (anonymous):

rewrite \(\sin(x)+\sqrt{3}\cos(x)\) as \(r\sin(x+\theta)\)

OpenStudy (anonymous):

\(r=\sqrt{1^2+\sqrt{3}^2}=\sqrt{4}=2\)

OpenStudy (anonymous):

\(\sin(\theta)=\frac{\sqrt{3}}{2}\) and \(\cos(\theta)=\frac{1}{2}\) so \(\theta=\frac{\pi}{3}\)